OpenStudy (anonymous):

A group of 8 friends (5 girls and 3 boys) plans to watch a movie, but they have only 5 tickets. How many different combinations of 5 friends could possibly receive the tickets?

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

\[_{nC _{r}}=\frac{ n! }{ (n-r)!r!}\] \[_{nC _{r}}=\frac{ 8! }{ (8-5)!5! }=56\] So there are 56 combinations.